amantha owns and operates a barber shop in the Houston area. The barber shop has 10 barber chairs. Samantha only operates 5 chairs at a time, leaving the rest of the chairs vacant to reduce overcrowding, clean-up time, and repair costs. Based on her latest revenue reports, she is averaging 3 customers receiving a haircut while the other two chairs are vacant. The barber shop is open 10 hours per day. Each customer requires 30 minutes to receive a proper haircut.

a) What is Samantha's daily design capacity?
b) What is Samantha's daily effective capacity?
c) What is Samantha's actual output per day?
d) What is Samantha's utilization rate?
e) Samantha wants to improve her utilization rate. Identify 3 ways she can do so.
Answer #1

Each customer requires 30 mins, that is 2 customer per hour

Working hours = 10 hours

Therefore, on 1 chair there can be 2*10= 20 customers per day

Answer a: Barber shop has 10 chairs , hence the design capacity = 10*20= 200 customers per day

Answer b: Effective chairs used is 5, hence the effective capacity = 5*20= 100 customers per day

Answer c: only average of 3 chairs are occupied at a time, actual output= 3*20= 60 customers per day

Answer d: Utilization= actual output/ design capacity = 60/200 = 0.3, 30%

Answer e: she can improve her utilization rate by

  1. Selling 5 chairs, this will reduce the design capacity and utilization will increase
  2. She can open new store and use extra 5 chairs there since she is not using it here in Houston area
  3. She can invest in marketing her shop to increase the customers footprint or by providing discount. More customers will increase the actual output of the shop